# Ultra-High-Energy Cosmic Rays from Neutrino-Emitting Tidal Disruption Events

### v2

@@ 6 Discussion
Choice of parameters, degeneracies. In this study, we fixed several key parameters such as the SMBH mass, the dissipation efficiency, the minimum energy and the spectral index of injected nuclei. We also took into account that certain parameters – including the SMBH mass and the IR, OUV, and X-ray luminosities – are constrained by observations. For instance, we fixed the non-thermal cosmic ray luminosity to be , where the mass accretion rate is taken to be proportional to the Eddington luminosity (and therefore to the SMBH mass). We then derived the local TDE rates that are required to power the observed UHECR flux. These parameters are, however, somewhat degenerate. Indeed, lower can be compensated by a higher local rate or a higher (higher SMBH mass). Similarly, a higher minimum nuclei energy or a harder injection spectrum lowers the energetic requirement as more non-thermal power is allocated at higher energies, see discussion in Winter (2024). Due to this degeneracy, our derived local rates should only be interpreted as indicative. Moreover, degeneracies should be taken into account when comparing with other works; for example, Biehl et al. (2018b) only fit the product of local rate and baryonic loading.
@@
Cosmological TDE rates, connection with jetted TDEs. Despite the existing degeneracies, it is interesting to compare the TDE rates found in this work with those expected from theory and observations. The theoretically expected local rate of all TDEs, (depending on the minimal SMBH mass; see Sec. 3.1) serves as an upper limit for the rate of the sub-class of UHECR-emitting TDEs. Our fit results are consistent with such limit: we found for the aalc-group and for the fdr-group, which amount to a fraction of the maximum value.
@@
It is noteworthy that our TDE fraction range is roughly compatible with the fraction of jetted TDEs, to (Piran & Beniamini, 2023), which suggests a possible connection with the existence of jets in TDEs. In the model of Winter & Lunardini (2023), the jet, which can be completely off-axis, may be the accelerator; the non-thermal UHECRs isotropize in the larger region enclosed by the dust torus, thus making the source effectively a quasi-isotropic emitter. On the other hand, a jet directly pointing towards us, such as in the model of Winter & Lunardini (2021), is expected to lead to significant electromagnetic signatures (e.g., non-thermal X-rays) which have not been observed in the three cases of neutrino-coincident TDEs but may account for the luminous X-ray afterglows of jetted TDEs, such as the recently reported AT 2022cmc (e.g., Andreoni et al., 2022; Pasham et al., 2023; Yuan et al., 2024b).
Earlier approaches, such as Biehl et al. (2018b), also used jetted TDEs for the combined description of UHECR and neutrino data. Note that from the UHECR perspective, there is no conceptual difference between those earlier models and our approach, as UHECRs are deflected by extragalactic magnetic fields and, therefore, do not directly trace the original direction of the jet. In fact, the value of the local (apparent) rate of jetted TDEs that was assumed in Biehl et al. (2018b), 0.1 , corresponds - after taking into account the beaming factor (opening angle from Piran & Beniamini (2023)) – to a local TDE rate 200 , which agrees very well with the values we found for the aalc-group.
@@
For neutrinos, however, a jetted model may lead to a higher flux at Earth, due to the contribution of neutrinos from internal shocks inside the jet, which can be very efficient. This additional flux comes from jets pointing in our direction, such as in Biehl et al. (2018b). Type of disrupted stars, nuclear composition. We found that the best-fit composition for the aalc-group is a mix of medium- and light-mass elements. It is overall similar to that of ONeMg-WD (in accordance with Zhang et al. (2017)), but requires additional light elements. The latter feature might be due to the loss of intermediate elements before they are transferred to the acceleration zone; it could also indicate the existence of two subgroups within the aalc-group, one involving white dwarfs and the other main-sequence stars. However, such a model is beyond the scope of our current analysis with the available data. For the fdr-group, the composition is likely dominated by iron or silicon, within the 1 AICc uncertainty region. Such a composition is also broadly consistent with an ONeMg-WDs.
@@
the previously mentioned local TDE rates, which are consistent with MS – especially for the aalc-group – and too high for consistency with WD (the TDE-WD rate is , see Sec. 3.1).
@@
The estimated black hole masses. In the current understanding of TDEs white dwarfs can only be disrupted by intermediate-mass black holes with typical masses around , while all three observed TDEs involve supermassive black holes around .
@@
There are potential solutions to this tension, where additional physical effects bring the required nuclear composition into consistency with the disruption of MS stars. A possibility is a strong enhancement of heavier elements through the injection processes. A simple, power-law enhancement factor might reproduce the required iron-like composition for the fdr-group (see Sec. 5), whereas for the aalc-group, a more complicated process that favors medium-mass elements would be required. Another idea is that during the disruption of a main-sequence star, central densities, and temperatures can dramatically increase for a very short period, leading to significant nuclear burning. However, there is no clear consensus on whether these conditions are sufficient: Luminet & Pichon (1989) argue in favor, while Guillochon et al. (2009) come to a negative answer. Thus, a self-consistent picture has not emerged yet. Such a picture will require a more comprehensive (UHECR-emitting) population model, which would not rely as much on the observed neutrino-TDE coincidences. Indeed, these may not be representative of the whole population; rather, they could be outliers with extraordinarily large neutrino fluences.
@@
During the completion of this work, a new study appeared (Milán Veres et al., 2024), revealing a re-brightening of AT2019aalc four years after its discovery. This new flare has characteristics – such as a dust echo and X-ray emission – that are similar to the first one. While the classification of AT2019aalc as TDE still remains unclear, it is speculated that its peculiar properties might be explained by the phenomenology of TDEs in AGN, for example, partial disruption. Spectroscopically, the flare source can be classified as Bowen Fluorescence Flare (BFF, a relatively new class of flaring AGN, Makrygianni et al., 2023). Interestingly, another such BFF with a very similar emission spectrum, AT2021loi, has been identified and has been found to be within the (rectangular) 90% confidence region of neutrino event IC-230511. This neutrino arrived 680 (290) days after the first (second) optical peak of AT2021loi.444In fact, an additional neutrino-flare pair was found, involving AT2017bgt and the neutrino IC-200410A, which was, however, a poorly reconstructed cascade event. These observations indicate that, indeed, AT2019aalc might be exceptional and that our aalc-group may correspond to a special type of TDEs in AGN related to BFFs, including AT2019aalc, AT2021loi, and, possibly, AT2017bgt. The spectroscopic results indicate that intermediate to heavy elements, required to describe the UHECR data, are indeed available in the system, even though a direct link to our TDE model requires more information on the origin of the emission lines, which is not yet available.
@@
Broader connections: individual extreme UHECR and neutrino events. Regarding the derived UHECR energies, the fdr-group might be able to produce particles with high energies, such as the eV particle observed in 1991 (Bird et al., 1995) and the Amaterasu particle with an energy of eV in 2021 (Abbasi et al., 2023). We note that Unger & Farrar (2024) suggests that a transient source may produce the Amaterasu particle, which positions TDEs as a potential candidate. Moreover, Bourriche & Capel (2024) proposed that the Amaterasu particle could have originated from emitted iron nuclei, which aligns with our results for the composition of the fdr-group. Recently, the detection of a high-energy neutrino event by KM3NeT-ARCA, with energies beyond tens of PeVs was reported Coelho (2024). This estimated energy is compatible with the energy where our best-fit spectra for the diffuse flux from both groups have a maximum, see Fig. 6.
@@
Fit quality and possible developments. The best value obtained in our analysis was approximately 2.5, which indicates that the fit quality is not ideal. This suggests that while our model captures the broad characteristics of the data, there are likely residuals and complexities in the data that are not fully accounted for. However, attempts to further improve the fit by introducing additional parameters or making the model more complex are unlikely to yield significant benefits. The reason is that the current data may not provide enough additional information to justify an increased model complexity, and such adjustments might only make the theoretical model overly intricate and challenging to interpret. Even in more straightforward cases, such as those discussed by Heinze et al. (2019) and Aab et al. (2017), achieving a perfect fit has proven challenging.
@@
A potential development of our analysis might be including information on the arrival direction of the observed UHECRs. Considering that this exercise has turned out to be difficult even with simple source models, as indicated by Capel & Mortlock (2019) and Abdul Halim et al. (2024a), we feel that applying it to our case of a realistic source model might introduce excessive complication while not adding significant improvements.
@@ 7 Summary and conclusions
For several years, TDEs have been discussed as potential emitters of UHECR and high-energy neutrinos. A significant advancement occurred recently, with the proposed associations between neutrinos detected at IceCube and the candidate TDEs AT2019dsg, AT2019fdr, and AT2019aalc. All of these events have two common features. The first is strong dust echoes, due to the absorption of optical-UV and X-ray photons by dust, and consequent re-emission in the IR band. The peak of the IR flare is observed as delayed relative to the blackbody peak, by the travel time from the central source to the dust region, days. Remarkably, using IR echoes as a selection criterion has led to the identification of AT2019aalc and its associated neutrino. The second commonality is that the coincident neutrinos have all been detected with similar delays of hundreds of days, which place them close to the dust echo peaks. These two facts suggest that the dust echoes may contribute significantly to neutrino production, by providing the target photons for hadron-photon scattering where neutrinos originate. Intriguingly, the primary nuclei must have energies in the UHECR range to exceed the photo-pion production threshold off the infrared targets, therefore the neutrino-TDE associations might be interpreted as indirect evidence for the acceleration of UHECRs in TDEs. In light of these associations, we have revisited the idea that TDEs may power the observed diffuse UHECR flux.
@@
We extended a quasi-isotropic neutrino production model, which was previously proposed to interpret the neutrino-TDE associations, by injecting a variety of nuclear species and allowing the relevant parameters (size of production region and maximal rigidity) to vary. We combined that with a UHECR transport code to predict the composition of cosmic rays at Earth. Finally, we performed a fit to the UHECR spectrum and composition data from Auger, and, for each set of fitted parameters, we computed the predicted diffuse neutrino flux at IceCube. It is important to note that the expected local rate of TDEs is determined by fitting the normalization (required emissivity) of the computed cosmic ray spectra to UHECR data, and that predicted diffuse neutrino fluxes are an output (rather than an input) of our calculation.
The main challenges for this work have been: (i) extrapolating from a few neutrino-emitting TDEs to a population of UHECR-emitting TDEs, (ii) reconciling the favored local TDE rate with the nuclear injection composition required by the data, and (iii) ensuring consistency with the individual neutrino-TDE associations. We have identified a best-fit scenario where two populations of TDEs, one being AT2019aalc-like and the other AT2019fdr-like, describe all relevant UHECR data best while meeting existing constraints.
@@
In such a scenario, the aalc-group has an injection composition that consists of a mix of light and intermediate (Na-Si) isotopes, and dominates the UHECR spectrum. In contrast, the fdr-group is characterized by a Fe-like composition and reproduces the highest energy tail of the spectrum. Therefore, our results indicate that at least two different classes of TDEs contribute to the observed UHECR flux. The inferred parameters (radius of production region and maximal rigidity) are found to be in consistency with earlier works. The expected local TDE rates range between 10 and 100 yr-1Gpc-3, where a significant uncertainty is allowed by degeneracies in the parameters. These rates correspond to a fraction of the overall TDE rate, which, interestingly, is compatible with the rate of jetted TDEs.
@@
From our results, it is not clear what type of stars must be disrupted to power the observed UHECRs. While white dwarf disruptions are favored from the injection composition perspective, main sequence stars are preferred from the local rate standpoint, and are supported by the estimated black hole and disrupted star masses of AT2019fdr, and AT2019aalc. As an option to resolve this discrepancy, we have proposed an enhancement of the injection of heavier elements in the acceleration process, which means that these elements should exhibit larger non-thermal contributions compared to the lighter nuclei. We found that the simplest realization of this idea could make the data compatible main sequence star disruptions, a more complicated enhancement mechanism that is most effective for intermediate-mass elements (rather than heavy elements) would be needed to provide a good fit. During completion of this work, evidence that AT2019aalc belongs to a special type of events (which can be identified with Bowen Fluorescence Flares – BFF) was presented, which might be related to (partial) TDEs in AGN. New candidate neutrino associations involving members of this group were identified. It is remarkable that our UHECR data analysis independently led to the conclusion that the aalc-group might be a special class, thus substantiating the idea further.
@@
The diffuse neutrino fluxes predicted by our best-fit model peak at 10-100 PeV energies, which are beyond the range of optimal sensitivity of the IceCube detector. Such peak energy suggests that a TDE could be the source of the recently observed high-energy neutrino by KM3NeT-ARCA. Future radio detection instruments like IceCube-Gen2 (black, Aartsen et al., 2019), RNO-G (olive, Aguilar et al., 2021), and GRAND200k (red, Álvarez Muñiz et al., 2019), should be able to robustly test our prediction for the neutrino flux.
@@
In conclusion, we find it overall plausible that TDEs could power the UHECRs in the light of recent neutrino-TDE associations. However, the extrapolation from a few neutrino-emitting TDEs to a population of UHECR-emitting TDEs is non-trivial due to limited knowledge and possible selection effects. Future data will potentially improve this situation, both on the TDE side and the neutrino side.

### v3

@@ 6 Discussion
In this study, we fixed several key parameters, such as the SMBH mass, the dissipation efficiency, the minimum energy, and the spectral index of injected nuclei. We also took into account that certain parameters – including the SMBH mass and the IR, OUV, and X-ray luminosities – are constrained by observations. For instance, we fixed the non-thermal cosmic ray luminosity to be , where the mass accretion rate is taken to be proportional to the Eddington luminosity (and therefore to the SMBH mass). We then derived the local TDE rates that are required to power the observed UHECR flux. These parameters are, however, somewhat degenerate. Indeed, lower can be compensated by a higher local rate or a higher (higher SMBH mass).
Similarly, a higher minimum nuclei energy or a harder injection spectrum lowers the energetic requirement as more non-thermal power is allocated at higher energies, see discussion in W. Winter (2024). Due to this degeneracy, our derived local rates should only be interpreted as indicative. Moreover, degeneracies should be taken into account when comparing with other works; for example,
D. Biehl et al. (2018b) only fit the product of local rate and baryonic loading.
@@
Cosmological TDE rates, connection with jetted TDEs. Despite the existing degeneracies, it is interesting to compare the TDE rates found in this work with those expected from theory and observations. The theoretically expected local rate of all TDEs, (depending on the minimal SMBH mass; see Sec. 3.1) serves as an upper limit for the rate of the sub-class of UHECR-emitting TDEs. Our fit results are consistent with such limit: we found yr-1Gpc-3 for the aalc-group and yr-1Gpc-3 for the fdr-group, which amount to a fraction of the maximum value.
@@
It is noteworthy that our TDE fraction range is roughly compatible with the fraction of jetted TDEs, to (T. Piran & P. Beniamini, 2023), which suggests a possible connection with the existence of jets in TDEs. In the model of W. Winter & C. Lunardini (2023), the jet, which can be completely off-axis, may be the accelerator; the non-thermal UHECRs isotropize in the larger region enclosed by the dust torus, thus making the source effectively a quasi-isotropic emitter. On the other hand, a jet directly pointing towards us, such as in the model of W. Winter & C. Lunardini (2021), is expected to lead to significant electromagnetic signatures (e.g., non-thermal X-rays) which have not been observed in the three cases of neutrino-coincident TDEs but may account for the luminous X-ray afterglows of jetted TDEs, such as the recently reported AT 2022cmc (e.g., I. Andreoni et al., 2022; D. R. Pasham et al., 2023; C. Yuan et al., 2024c, b). Earlier approaches, such as D. Biehl et al. (2018b), also used jetted TDEs for the combined description of UHECR and neutrino data. Note that from the UHECR perspective, there is no conceptual difference between those earlier models and our approach, as UHECRs are deflected by extragalactic magnetic fields and, therefore, do not directly trace the original direction of the jet. In fact, the value of the local (apparent) rate of jetted TDEs that was assumed in D. Biehl et al. (2018b), 0.1 yr-1Gpc-3, corresponds - after taking into account the beaming factor (opening angle from T. Piran & P. Beniamini (2023)) – to a local TDE rate 200 yr-1Gpc-3, which agrees very well with the values we found for the aalc-group.
For neutrinos, however, a jetted model may lead to a higher flux at Earth, due to the contribution of neutrinos from internal shocks inside the jet, which can be very efficient. This additional flux comes from jets pointing in our direction, such as in D. Biehl et al. (2018b).
@@
Type of disrupted stars, nuclear composition. We found that the best-fit composition for the aalc-group is a mix of medium- and light-mass elements. It is overall similar to that of ONeMg-WD (in accordance with B. T. Zhang et al. (2017)), but requires additional light elements. The latter feature might be due to the loss of intermediate elements before they are transferred to the acceleration zone; it could also indicate the existence of two subgroups within the aalc-group, one involving white dwarfs and the other main-sequence stars. However, such a model is beyond the scope of our current analysis with the available data. For the fdr-group, the composition is likely dominated by iron or silicon, within the 1 AICc uncertainty region. Such a composition is also broadly consistent with an ONeMg-WDs.
@@
The local TDE rates. Our required rates for the aalc-group are well consistent with MS stars, but somewhat in tension with the predicted rates of disruption of WD stars (unless the most extreme values of TDE-WD rates are invoked, see Sec. 3.1).
@@
The estimated black hole masses. In the current understanding of TDEs white dwarfs can only be disrupted by intermediate-mass black holes with typical masses around , while all three observed TDEs involve supermassive black holes around . Based on that, Disruption of the WD is unlikely for AT2019dsg, AT2019fdr, and AT2019aalc. Therefore, for these TDEs, the MS star is a more probable candidate.
@@
There are potential solutions to this tension, where additional physical effects bring the required nuclear composition into consistency with the disruption of MS stars. A possibility is a strong enhancement of heavier elements through the injection processes. A simple, power-law enhancement factor might reproduce the required iron-like composition for the fdr-group (see Sec. 5), whereas for the aalc-group, a more complicated process that favors medium-mass elements would be required. Another idea is that during the disruption of a main-sequence star, central densities, and temperatures can dramatically increase for a very short period, leading to significant nuclear burning. However, there is no clear consensus on whether these conditions are sufficient: J. P. Luminet & B. Pichon (1989) argue in favor, while J. Guillochon et al. (2009) come to a negative answer. Thus, a self-consistent picture has not emerged yet. Such a picture will require a more comprehensive (UHECR-emitting) population model, which would not rely as much on the observed neutrino-TDE coincidences. Indeed, these may not be representative of the whole population; rather, they could be outliers with extraordinarily large neutrino fluences.
@@
During the completion of this work, a new study appeared (P. Milán Veres et al., 2024), revealing a re-brightening of AT2019aalc four years after its discovery. This new flare has characteristics – such as a dust echo and X-ray emission – that are similar to the first one. While the classification of AT2019aalc as TDE still remains unclear, it is speculated that its peculiar properties might be explained by the phenomenology of TDEs in AGN, for example, partial disruption. Spectroscopically, the flare source can be classified as Bowen Fluorescence Flare (BFF, a relatively new class of flaring AGN, L. Makrygianni et al., 2023). Interestingly, another such BFF with a very similar emission spectrum, AT2021loi, has been identified and has been found to be within the (rectangular) 90% confidence region of neutrino event IC-230511. This neutrino arrived 680 (290) days after the first (second) optical peak of AT2021loi.555In fact, an additional neutrino-flare pair was found, involving AT2017bgt and the neutrino IC-200410A, which was, however, a poorly reconstructed cascade event. These observations indicate that, indeed, AT2019aalc might be exceptional and that our aalc-group may correspond to a special type of TDEs in AGN related to BFFs, including AT2019aalc, AT2021loi, and, possibly, AT2017bgt. The spectroscopic results indicate that intermediate to heavy elements, required to describe the UHECR data, are indeed available in the system, even though a direct link to our TDE model requires more information on the origin of the emission lines, which is not yet available. Nevertheless, the limited sample size, classification ambiguities, and potential AGN contamination warrant caution against drawing strong conclusions from these individual events.
Compatibility between observed neutrino energies and predicted neutrino spectra. The direct connection between neutrino energy and CR rigidity (cf., discussion after Eq. (1), applied to nuclei heavier than hydrogen) means that neutrinos are produced at energies following – much higher than the inferred reconstructed neutrino energies in the 100 TeV range for the three neutrino-TDE associations. A quantification of the statistical compatibility between our predicted neutrino spectra (see e.g. Fig. 6, upper right panel) and the energy uncertainty of the IceCube events is nevertheless not possible because uncertainties for the three TDE-associated events have not been published so far, and (especially for muon tracks) the energy estimate is strongly event-dependent and non-Gaussian; it depends e.g. on the uncertainty on the vertex distance and how much energy was lost before reaching the detector (which means it depends on the direction, depth, etc.).
An additional challenge is that the detector response depends on the spectral shape itself. This can be seen, for instance, in the neutrino association of IC-170922A with the flaring AGN blazar TXS 0506+056 for a similar event, where the most probable neutrino energy was given to be 290 TeV: the harder the assumed spectrum in that energy range, the more the probability distribution prefers higher reconstructed neutrino energies (IceCube Collaboration, 2013, see Suppl. Materials, Fig. S2); our predicted spectrum is even much harder in that energy range. A generic attempt to quantify these uncertainties for power law spectra has been made in W. Winter & C. Lunardini (2023) (see e.g. gray-shaded areas in Fig. 10).
This problem has been, in fact, studied in greater detail for AGN blazars in X. Rodrigues et al. (2024), where the predicted neutrino energy follows the maximal proton energy – which is frequently used as free parameter in AGN blazar models: if it is chosen to be in the few-PeV range, it can produce neutrino spectra peaking in the 100 TeV energy range. At these energies, the target photon number is however low, which means that super-Eddington proton luminosities need to be invoked to produce reasonable neutrino event rates, which are incompatible with the standard accretion paradigm in AGN. A higher proton energy can significantly reduce this tension, which leads to the hypothesis that AGN blazar neutrino spectra may also peak at higher energies. X. Rodrigues et al. (2024) have therefore analyzed the detector response for these spectral cases in greater detail (see App. A) and have come to the come to the conclusion that for -like spectra much higher neutrino energies are compatible with the detector response (see Fig. A.2). In fact, the predicted spectra in X. Rodrigues et al. (2024) therefore peak at around the same energies as our TDE predicted spectra. While the theoretical arguments are different here, the conclusions on the detector response apply similarly. However, in F. Oikonomou et al. (2021) the probability that a 20 PeV neutrino is reconstructed with a muon energy as low as 210 TeV was estimated for the individual event IC-190730A associated with an AGN blazar to be around 25%. While such an individual event is consistent with the hypothesis that the spectrum peaks at higher energies, it raises the question if three independent events from three different astrophysical objects with such a spectrum can be found at such low energies, which may be at the level of following their arguments. This line of argumentation would, however, also apply to set of AGN blazars, which raises two more generally applicable possibilities: 1) The observed neutrinos (with the exception of KM3-230213A perhaps, see below) are not coming from a population of sources producing the UHECRs or 2) Some systematics on the detector side leads to an offset in the reconstructed neutrino energies.
Another hint for hard TDE spectra comes from stacking searches, where too soft spectra limit the possibility of neutrino-TDE associations (J. Necker, 2023).
@@
Broader connections: individual extreme UHECR and neutrino events. Regarding the derived UHECR energies, the fdr-group might be able to produce particles with high energies, such as the eV particle observed in 1991 (D. J. Bird et al., 1995) and the Amaterasu particle with an energy of eV in 2021 (R. U. Abbasi et al., 2023). We note that M. Unger & G. R. Farrar (2024) suggests that a transient source may produce the Amaterasu particle, which positions TDEs as a potential candidate. Moreover, N. Bourriche & F. Capel (2024) proposed that the Amaterasu particle could have originated from emitted iron nuclei, which aligns with our results for the composition of the fdr-group. Recently, the detection of a high-energy neutrino event KM3-230213A by KM3NeT-ARCA, with a median neutrino energy of 220 PeV (range 72 PeV to 2.6 EeV at the 90% CL) was reported (S. Aiello et al., 2025), with no clear source association. This energy range is compatible with our diffuse flux computation in Fig. 6, even though our spectrum peaks at slightly lower energies. It is interesting that the interpretation of this event coming from a transient with a lifetime comparable or shorter than the live time of the experiment (335 days) alleviates the tension between IceCube and KM3NeT data (S. W. Li et al., 2025; A. Neronov et al., 2025) – compared to other possible explanations, such as cosmogenic neutrinos.
@@
Fit quality and possible developments. The best value obtained in our analysis was approximately 2.5, which indicates that the fit quality is not ideal. This suggests that while our model captures the broad characteristics of the data, there are likely residuals and complexities in the data that are not fully accounted for. However, attempts to further improve the fit by introducing additional parameters or making the model more complex are unlikely to yield significant benefits. The reason is that the current data may not provide enough additional information to justify an increased model complexity, and such adjustments might only make the theoretical model overly intricate and challenging to interpret. Even in more straightforward cases, such as those discussed by J. Heinze et al. (2019) and A. Aab et al. (2017), achieving a perfect fit has proven challenging.
@@
A potential development of our analysis might be including information on the arrival direction of the observed UHECRs. Considering that this exercise has turned out to be difficult even with simple source models, as indicated by F. Capel & D. J. Mortlock (2019) and A. Abdul Halim et al. (2024a), we feel that applying it to our case of a realistic source model might introduce excessive complication while not adding significant improvements.
@@ 7 Summary and conclusions
For several years, TDEs have been discussed as potential emitters of UHECR and high energy neutrinos. A significant advancement occurred recently, with the proposed associations between neutrinos detected at IceCube and the TDEs/candidates, AT2019dsg, AT2019fdr, and AT2019aalc. All of these events have two common features. The first commonality are strong dust echoes, due to the absorption of optical-UV and X-ray photons by dust and consequent re-emission in the IR band. The peak of the IR flare is observed as delayed relative to the blackbody peak, by the travel time from the central source to the dust region, days. Remarkably, using IR echoes as a selection criterion has led to the identification of AT2019aalc and its associated neutrino. The second commonality is that the coincident neutrinos have all been detected with similar delays of hundreds of days, which place them close to the dust echo peaks. These two facts suggest that the dust echoes may contribute significantly to neutrino production, by providing the target photons for hadron-photon scattering where neutrinos originate. Intriguingly, the primary nuclei must have energies in the UHECR range to exceed the photo-pion production threshold off the infrared targets, therefore the neutrino-TDE associations might be interpreted as indirect evidence for the acceleration of UHECRs in TDEs. In light of these associations, we have revisited the idea that TDEs may power the observed diffuse UHECR flux.
@@
We extended a quasi-isotropic neutrino production model, which was previously proposed to interpret the neutrino-TDE associations, by injecting a variety of nuclear species and allowing the relevant parameters (size of production region and maximal rigidity) to vary. We combined that with a UHECR transport code to predict the cosmic ray composition at Earth. Finally, we performed a fit to the UHECR spectrum and composition data from Auger, and, for each set of fitted parameters, we computed the predicted diffuse neutrino flux at IceCube. It is important to note that the expected local rate of TDEs is determined by fitting the normalization (required emissivity) of the computed cosmic ray spectra to UHECR data, and that predicted diffuse neutrino fluxes are an output (rather than an input) of our calculation.
@@
The main challenges for this work have been: (i) extrapolating from a few neutrino-emitting TDEs to a population of UHECR-emitting TDEs, (ii) reconciling the favored local TDE rate with the nuclear injection composition required by the data, and (iii) ensuring consistency with the individual neutrino-TDE associations. We have identified a best-fit scenario where two populations of TDEs, one being AT2019aalc-like and the other AT2019fdr-like, that describe all relevant UHECR data best while meeting existing constraints. In such a scenario, the aalc-group has an injection composition that consists of a mix of light and intermediate (Na-Si) isotopes, and dominates the UHECR spectrum. In contrast, the fdr-group is characterized by a Fe-like composition and reproduces the highest energy tail of the spectrum. Therefore, our results indicate that at least two different classes of TDEs contribute to the observed UHECR flux. The inferred parameters (radius of production region and maximal rigidity) are found to be in consistency with earlier works. The expected local TDE rates range between 10 and 100 yr-1Gpc-3, where a significant uncertainty is allowed by degeneracies in the parameters. These rates correspond to a fraction of the overall TDE rate, which, interestingly, is compatible with the fraction of jetted TDEs.
@@
From our results, it is not clear what type of stars must be disrupted to power the observed UHECRs. White dwarf disruptions seem to have the right mass composition, but are too rare and associated with lighter black hole masses. Main sequence stars are preferred as progenitors because their local TDE rate is high enough, and the estimated black hole masses and star’s masses of AT2019fdr, and AT2019aalc favor this hypothesis. However, their composition (in terms of the mass fraction) is light. As an option to resolve this discrepancy, we have proposed an enhancement of the injection of heavier elements in the acceleration process, which means that these elements should exhibit larger non-thermal contributions compared to the lighter nuclei. We found that the simplest realization of this idea could make the data compatible with main sequence star disruptions, but a more complicated enhancement mechanism that is most effective for intermediate-mass elements (rather than heavy elements) would be needed to provide a good fit. During completion of this work, evidence that AT2019aalc belongs to a special type of events (which can be identified with Bowen Fluorescence Flares – BFF) was presented, which might be related to (partial) TDEs in AGN. New candidate neutrino associations involving members of this group were identified. It is remarkable that our UHECR data analysis leads to the same conclusion, that the aalc-group might be a special class, thus substantiating the idea of two distinct stellar populations or classes of events further.
In order to associate a specific astrophysical class with UHECRs, reproducing the observed spectrum and composition at Earth is a necessary condition. However, direct source identification remains challenging, as UHECRs are not only significantly deflected by intergalactic magnetic fields but also experience long propagation time delays (R. Mbarek & D. Caprioli, 2025). These delays make it impossible to directly associate UHECRs with transient sources like TDEs. Instead, source and propagation simulations provide a valuable approach to investigating whether certain source classes, such as TDEs, can reproduce the observed UHECR spectrum and composition. At the same time, these simulations help identify key challenges and constraints on the physical conditions required for TDEs to be viable UHECR sources.
We note that the predicted neutrino flux associated with UHECRs inevitably peaks at 10-100 PeV, whereas the observed neutrino energies had much lower inferred energies in the 100 TeV range; see Sec. 6 for a discussion of the compatibility.
@@
On the other hand, such high peak energies suggests that a TDE or a TDE-like AGN flare (C. Yuan et al., 2025) could be the source of the recently observed ultra-high-energy neutrino by KM3NeT-ARCA (S. Aiello et al., 2025). Future radio detection instruments like IceCube-Gen2 (black, M. G. Aartsen et al., 2019), RNO-G (olive, J. Aguilar et al., 2021), and GRAND200k (red, J. Álvarez Muñiz et al., 2019), should be able to robustly test our prediction for the neutrino flux.
@@
In conclusion, we find it overall plausible that TDEs could power the UHECRs in the light of recent neutrino-TDE associations. However, the extrapolation from a few neutrino-emitting TDEs, which may not be archetypic and may be also not very statistically confident in terms of their directionality, to a population of UHECR-emitting TDEs is non-trivial due to limited knowledge and possible selection effects – and some of the parameters, such as the observed neutrino energies, do not support this hypothesis. Independent of that, TDEs are interesting UHECR candidates for different reasons, such as their transient nature, high enough emissivity, and negative source evolution, and our UHECR model and analysis (including the diffuse neutrino flux prediction) remain valid independent of the neutrino associations. Future data will potentially improve the predictions, both on the TDE side and the neutrino side.